# Toward A Causal Framework for Modeling Perception

### v3

@@ Abstract
Perception occurs when individuals interpret the same information differently. It is a known cognitive phenomenon with implications for bias in human decision-making. Perception, however, remains understudied in machine learning (ML). This is problematic as modern decision flows, whether partially or fully automated by ML applications, always involve human experts. How might we account for cases in which two experts, e.g., interpret differently the same deferred instance or explanation from a ML model? Addressing this and similar questions requires a formulation of perception, particularly, in a manner that integrates with ML-enabled decision flows. In this work, we present a first approach to modeling perception causally. We define perception under causal reasoning using structural causal models (SCM). Our approach formalizes individual experience as additional causal knowledge that comes with and is used by the expert decision-maker in the form of a SCM. We define two kinds of probabilistic causal perception: structural perception and parametrical perception. We showcase our framework through a series of examples of modern decision flows. We also emphasize the importance of addressing perception in fair ML, discussing relevant fairness implications and possible applications.
@@ 1 Introduction
The same piece of information can be interpreted differently by individuals. The late Kahneman (2021) refers to this cognitive phenomenon as perception. It is a product of the mental heuristics humans use to enable fast decision-making under uncertainty. These heuristics and, in turn, perception are shaped by individual experience, which can lead to biased decision-making. The Linda Problem (Tversky and Kahneman 1974, 1981, 1983), for instance, illustrates how perception can override logical reasoning when individuals rely on intuitive judgment. In a series of famous experiments, participants often committed the conjunction fallacy by judging the conjunction (“Linda is a bank teller and a feminist”) as more probable than a single constituent (“Linda is a bank teller”) violating the basic laws of probability that govern rational decision-making. These and similar experiments (Kahneman 2011; Kahneman et al. 2016; Thaler and Sunstein 2008), along with the growing regulatory emphasis on fairness, transparency, and safety in machine learning (ML) applications (Álvarez et al. 2024), highlight the need to consider perception as a core problem in ML.
Modern decision processe are powered by ML applications and involve some degree of human decision-making (Ruggieri et al. 2023; Scantamburlo, Baumann, and Heitz 2024; Ruggieri and Pugnana 2025). In learning to defer (LtD), e.g., the goal is to learn a ML model that abstains from predicting on instances it is not certain of and defers the decision to a human expert (Madras, Pitassi, and Zemel 2018). In explainable artificial intelligence (xAI), e.g., the goal is to develop methods that explain to an expert the predictions of a ML model (Guidotti et al. 2019). Perception clearly can occur in both settings when multiple individual experts are involved, potentially impacting the outputs of these ML applications. How might we develop LtD and xAI methods when experts may disagree on the same deferred instance or explanation, respectively? These questions are often taken for granted when building modern decision flows (for a discussion, see Cabitza, Campagner, and Basile (2023); Fahimi et al. (2024); Srivastava, Heidari, and Krause (2019)). Addressing this gap requires a formalization of perception that aligns with standard ML problem formulations, which is probabilistic (Pearl 1988, 2009; Goodfellow et al. 2016). In this work, we provide a first approach to this objective.
@@ 1.1 Our Contributions
How can we formally represent individuals who disagree in their interpretation of the same information? We first propose a probabilistic framework for perception, in which individuals construct disagreeing probability distributions, $P(\mathbf{X})$, for the same information $\mathbf{X}$. Under this framework, we can explain, for example, why one expert assigns a higher probability to a positive classification for a deferred instance than another expert does. We then propose causal perception by extending the formulation to account for probabilistic causal reasoning. Here, individuals are equipped with structural causal models (SCM) (Pearl 2009) that not only describe $P(\mathbf{X})$ but also govern how they reason about hypothetical or counterfactual scenarios. As a result, individuals may disagree not only on observed distributions but also on interventional and counterfactual ones. Under this framework, we can describe how the two experts arrive at different classification probabilities by answering counterfactual questions on the deferred instance. Our framework draws on foundational work in probabilistic and causal reasoning under uncertainty (Pearl 1988, 2009). We adopt a particular view of perception as disagreement in the interpretation of information conveyed through probabilities.
We define two kinds of causal perception, structural and parametrical. In structural perception individuals disagree on the cause-effect relationships, while in parametrical perception individual disagree on the causal effect between an established cause-effect pair. We complement these two kinds of causal perception by providing an initial approach to modeling the processes by which individuals construct their own SCMs. Further, we discuss areas of interest within fair ML and point at promising directions of future work involving perception, such as situated bias and loaded attributes. To showcase our framework, we present a series of examples involving ML applications and individual experts.
In the rest of the paper, we introduce the necessary background in Section 2. We then define probabilistic perception, or perception, in Section 3 and causal probabilistic perception, or causal perception, in Section 4. We present the intuition behind constructing the SCM in Section 5. We discuss fair ML implications in Section 6. We conclude in Section 7.
@@ 1.2 Related Work
Perception has been studied mainly by psychologist and views on perception vary within this field. In this work, broadly, we view perception as a difference in interpretation of the same information between individuals. Further, these multiple interpretations are driven by mental heuristics and materialize in the form of probabilities. The focus has been on the representativeness heuristic, where an event is made to be more representative of a class than what it actually is, as measured by a higher probability (Tversky and Kahneman 1974, 1981, 1983). We come back to these two points in Section 3. Such view falls within a more cognitive-bias-driven interpretation, in which perception influences rational decision-making under uncertainty (Kahneman 2011). There are, for instance, more constructivist (Gregory 1970) and sensorial (Goldstein 1996) views that focus on the interpretative aspect of perception. Within this view of perception, Bayesian modeling—in which the baseline representativeness of an event (the prior) is adjusted by the agent based on her experience (the posterior)—remains the common approach for modeling explicitly the representativeness heuristic and, thus, implicitly perception (Costello 2009; Bordalo et al. 2016; Tentori 2022). We are the first to use SCM under this view of perception, which allows to frame it also in terms of hypothetical distributions that represent counterfactual reasoning. Our work also extends previous works since SCM, or causal Bayesian networks (Pearl 2009), include Bayesian reasoning.
@@
ML researchers are increasingly interested in cognitive biases (e.g., Bengio (2019); Booch et al. (2021)). Such works focus on how to create ML-driven intelligent systems that improve over and potentially replace the irrational human expert. It is an active and interdisciplinary line of research (e.g., Agrawal, Gans, and Goldfarb (2019)). Our work adds to this growing line of research, making use of causal reasoning as a basis for modeling human-like reasoning (Schölkopf 2022; Pearl 2009). Further, there are narrower lines of ML work that study these cognitive biases and their impact on the human user of the ML application. We highlight those studying the human-in-the-loop problem from a fairness and accuracy perspective (e.g., De et al. (2020); Mozannar et al. (2023); Palomba et al. (2024)), which includes LtD (Madras, Pitassi, and Zemel 2018) and xAI (Rong et al. 2024). Such works consider the human’s interaction with the ML model, incorporating it within the problem formulation. With some exceptions (Caraban and Karapanos 2020; Rastogi et al. 2022; Yang, Folke, and Shafto 2023), these works, however, do not give agency to the human user and formulate it as an additional and costly decision-maker, ignoring any influence from individual experience. Different from these works, we explore the setting of having multiple users that disagree on the information provided. In that sense, our work gives more agency to the user by recognizing the role of perception among a set of heterogeneous users.
@@
Within fair ML, the role of these cognitive biases, with some exceptions (Echterhoff, Yarmand, and McAuley 2022; Bertrand et al. 2022) is largely unexplored. Some of these works show how the cognitive biases can be exacerbated by the xAI techniques used by the expert to explain a model (Bertrand et al. 2022) and can affect the expert’s evaluation of a model’s output (Echterhoff, Yarmand, and McAuley 2022). Given the problem of context-aware fairness, where we recognize that fairness has different meanings across humans, works like Srivastava, Heidari, and Krause (2019) and Yaghini, Krause, and Heidari (2021) design user experiments to test for the human perception of fairness. These experiments show that the fairness of an outcome can be judged differently depending on who is the individual judging. We add to these works by formalizing perception itself through causal reasoning. In Section 5.3 we further discuss perception within causal fair ML.
@@ 1.3 A Running Example
To illustrate our work, throughout the paper we use a series of ML-enabled decision flow scenarios based on the example below. Example 1 is an extended version of Kleinberg et al. (2019)’s college admissions tiebreaker example.
@@
(College Admissions) An admissions officer (the Decision Maker, or DM) chooses the incoming class based on the applicants’ profiles. Assume a decision flow in which the officer admits, $Y=1$, or rejects, $Y=0$, applicants based on their SAT results, $X_{1}$, high-school GPA, $X_{2}$, and suitability scores $f(X_{1},X_{2})=G\in[0,1]$ where $f$ is a ML model trained by the college. The officer, through the applicants’ motivation letters, also has access to their address, $Z$. Consider three scenarios in which the officer relies on $f$ with varying degrees for the decision:
@@
Scenarios (i) and (ii) represent a partially automated decision flow in which $f$ aids the DM, while scenario (iii) represents a fully automated decision flow in which $f$ replaces the DM. These are high-level but common ML-enabled scenarios. Additional context is provided for each scenario as we extend Example 1.
@@ 2 Background
We represent information as a set of $p$ discrete random variables $\mathbf{X}=X_{1},\dots,X_{p}$. Let $P(\mathbf{X})$ denote the joint probability distribution of $\mathbf{X}$ with $P(\mathbf{x})$ representing the joint probability that $\mathbf{X}$ equals the $p$ instances $\mathbf{x}=x_{1},\dots,x_{p}$. We treat individuals, like a human user, and objects, like a ML application, as agents.
@@
Borrowing from signaling games (Spence 1973), we consider two types of agents. A sender $S\in\mathcal{S}$, with $\mathcal{S}$ denoting the set of senders, is an agent that provides information while a receiver $R\in\mathcal{R}$, with $\mathcal{R}$ denoting the set of receivers, is an agent that interprets the provided information. These agents, however, are non-strategic and we consider a single one-shot game (Sobel 2020). In short, in our setup the sender sends and the receiver receives information without tricking one another.
@@
For readability, we focus on the additive noise models (ANMs), where $f_{j}$ is a linear transformation:
@@
with $\beta_{ij}\in\mathbb{R}$ denoting the causal weight of the $i$-th parent $pa(j)_{i}$ of $X_{j}$. Our discussion applies to all SCM, but we focus on the ANM case for illustrative purposes.
@@
which simplifies reasoning about $P(\mathbf{X})$, as it states that $X_{i}$ is conditionally independent of all other variables given its parents $X_{pa(i)}$. This property is known as the Markovian condition (Peters, Janzing, and Schölkopf 2017).
For the hypothetical scenarios, or what if, it is possible to generate new distributions of $P(\mathbf{X})$ by intervening the SCM $\mathcal{M}$. This is because SCMs build on an interventionist account of causality (Woodward 2005). An intervention on a single variable $X_{i}$ is done via the do-operator, $do(X_{i}=x_{i})$, which replaces the structural equation in $\mathbf{F}$ for the variable $X_{i}$ with the value $x_{i}$. Interventions apply similarly for multiple variables, $do(X_{i}=x_{i},X_{j}=x_{j})$, replacing the structural equations for each variable individually. Let $\mathcal{I}_{\mathbf{X}}$ denote the set of all interventions, which is an index set with each index representing a specific intervention on the variables $\mathbf{X}$. We use $\emptyset\in\mathcal{I}_{\mathbf{X}}$ to denote the null intervention. As Rubenstein et al. (2017) point out, $\mathcal{I}_{\mathbf{X}}$ has a natural partial ordering, in which for interventions $i,j\in\mathcal{I}_{\mathbf{X}},\,i\leq_{\mathbf{X}}j$ if and only if $i$ intervenes on a subset of the variables that $j$ intervenes on and sets them equal to the same values as $j$. For instance, $do(X_{i}=x_{i})\leq_{\mathbf{X}}do(X_{i}=x_{i},X_{j}=x_{j})$. Intuitively, it means the $j$ intervention can be done after the $i$ intervention without needing to change the modifications done by the $i$ intervention on the SCM $\mathcal{M}$.
@@
By definition, $P(\mathbf{\mathbf{X}})\in\mathcal{P}_{\mathbf{X}}$. Further, $\mathcal{P}_{\mathbf{X}}$ is a singleton comprised of $P(\mathbf{X})$ when $\mathcal{I}_{\mathbf{X}}=\{\emptyset\}$. Intuitively, $\mathcal{P}_{\mathbf{X}}$ represents all possible ways of reasoning causally about variables $\mathbf{X}$ as implied by a SCM $\mathcal{M}$.
@@ 3 Perception of Probability
We present perception in its most basic form. The goal is to formulate when two individuals, the receivers, interpret differently the same information from a ML application, formalized as the sender. We consider, in particular, the setting in which these individuals act as some sort of decision maker within a modern decision flow.
We deal with perception of probability as we assume that the individuals manifest their interpretations of the information received using probabilities. This assumption is rooted in how humans reason under uncertainty, which has a long and established tradition across multiple fields that study bias in human decision-making (see, e.g., Kahneman (2011); Thaler and Sunstein (2008)). We find this to be a reasonable and realistic assumption. Whether it is because the setting is inherently stochastic or the individual cannot possibly observe all the variables involved, there is always some degree of uncertainty involved in modern decision flows (Pearl 1988). Probabilistic reasoning helps capture the uncertainty behind making a choice based on a ML application’s outcome. Further, probability theory is one of the pillars of ML (Goodfellow et al. 2016). By describing perception in terms of probabilities, we can relate it to standard ML problem formulations. Let us now offer a preliminary definition of perception of probability entirely based on this intuition.
@@
(Perception) For receivers $R_{i},R_{j}\in\mathcal{R}$, given the information by sender $S\in\mathcal{S}$ in the form of a random variable $X$, perception occurs when, for a threshold ϵ+italic-$\epsilon\in\mathbb{R}^{+}$, we have that:
$d\big{(}P_{R_{i}}(X),P_{R_{j}}(X)\big{)}>\epsilon$ (5)
@@
(From variable to variables) Definition 3.1 naturally applies to the $p$ random variables $\mathbf{X}$. In that case, $P_{R_{i}}(\mathbf{X})$ and $P_{R_{j}}(\mathbf{X})$ represent the joint probability distributions of the receivers for information $\mathbf{X}$. Further, in such a setting, we can consider the particular case in which one of the random variables denotes the target $Y$ and $P_{R_{i}}(Y|\mathbf{X})$ and $P_{R_{j}}(Y|\mathbf{X})$ represent the conditional joint probability distributions of the receivers for information $\mathbf{X}$.
@@
(Instance-based perception) Implicit to Definition 3.1 is a disagreement at the instance level $X=x$, meaning $d\big{(}P_{R_{i}}(x),P_{R_{j}}(x)\big{)}>\epsilon$ where $P_{R_{i}}(x)$ and $P_{R_{i}}(x)$ are probabilities and $d(\cdot,\cdot)$ a suitable distance metric. [^2] Focus is on the specific instance as, indeed, distance among probability distributions in Equation 8 may be $\leq\epsilon$, with still a number of instances having distance $>\epsilon$. Intuitively, if a receiver can construct the probability distribution $P(X)$, then it can also compute the probability of any instance $x\in X$. It is also the case for the joint probabilities of the $p$ instances in $\mathbf{x}$ of the $p$ random variables in $\mathbf{X}$.
With Definition 3.1, we are interested in capturing how perception—as a disagreement in probability—would look like formally. We take for granted the process behind how these receivers construct(ed) $P(X)$ and focus on the fact that there is a (probabilistic) disagreement between them. Further, the fact that there is disagreement of probability between $R_{i}$ and $R_{j}$ only has an impact when these two receivers must make a decision within the same decision-making context based on $P(X)$. If $d\big{(}P_{R_{i}}(X),P_{R_{j}}(X)\big{)}\leq\epsilon$, then we would not have a disagreement and would be indifferent between $R_{i}$ and $R_{j}$. In that case, these two receivers, for our purposes, would be indistinguishable within said context. Let us consider the next example.
@@
Let us consider Example 1, scenario (i). Suppose the admissions officers $R_{1}$ and $R_{2}$ receive the same deferred applicant with profile $\mathbf{x}=\langle x_{1},x_{2}\rangle$ by the abstaining ML model $f$. Each officer constructs different $P(Y|X_{1},X_{2})$, e.g., with $P_{R_{1}}(Y|x_{1},x_{2})>P_{R_{2}}(Y|x_{1},x_{2})$ for which officer $R_{1}$ accepts while $R_{2}$ rejects the applicant.
As illustrated in Example 2, Definition 3.1 dwells with the observational probability distribution or what is. In other words, we do not allow for hypothetical scenarios: the applicant is judged based on what is observed in his or her application. There is no, for instance, counterfactual reasoning involved when deciding over the deferred applicant. We expand such a setting in Section 4.

### v4

@@ Abstract
Perception occurs when individuals interpret the same information differently. It is a known cognitive phenomenon with implications for bias in human decision-making. Perception, however, remains understudied in machine learning (ML). This is problematic as modern decision flows, whether partially or fully automated by ML applications, always involve human experts. For instance, how might we account for cases in which two experts interpret differently the same deferred instance or explanation from a ML model? Addressing this and similar questions requires first a formulation of perception, particularly, in a manner that integrates with ML-enabled decision flows. In this work, we present a first approach to modeling perception causally. We define perception under causal reasoning using structural causal models (SCMs). Our approach formalizes individual experience as additional causal knowledge that comes with and is used by the expert decision-maker in the form of a SCM. We define two kinds of probabilistic causal perception: structural and parametrical. We showcase our framework through a series of examples of modern decision flows. We also emphasize the importance of addressing perception in fair ML, discussing relevant fairness implications and possible applications.
@@ 1 Introduction
The same piece of information can be interpreted differently by individuals. The late Kahneman (2021) refers to this cognitive phenomenon as perception. It is a product of the mental heuristics humans use to enable fast decision-making under uncertainty. These heuristics and, in turn, perception are shaped by individual experience, which can lead to biased decision-making. The Linda Problem (Tversky and Kahneman 1974, 1981, 1983), for instance, illustrates how perception overrides logical reasoning when individuals rely on intuitive judgment. In a series of famous experiments, participants committed the conjunction fallacy by judging the conjunction (“Linda is a bank teller and a feminist”) as more probable than a single constituent (“Linda is a bank teller”), violating the basic laws of probability that govern rational decision-making. These and similar decision-making experiments (Kahneman 2011; Kahneman et al. 2016; Thaler and Sunstein 2008), together with the growing regulatory emphasis on trustworthy machine learning (ML) applications (Álvarez et al. 2024), underscore the need to consider perception as a core problem in ML.
Modern decision processes are powered by ML applications and always involve some degree of human decision-making (Ruggieri et al. 2023; Scantamburlo, Baumann, and Heitz 2024; Ruggieri and Pugnana 2025). In learning to defer (LtD), for example, the goal is to learn a ML model that abstains from predicting on instances it is not certain of and defers the decision to a human expert (Madras, Pitassi, and Zemel 2018). In explainable artificial intelligence (xAI), for example, the goal is to develop methods that explain to an expert the predictions of a ML model (Guidotti et al. 2019). Perception can clearly occur in both settings when multiple individual experts are involved, potentially impacting the outputs of these ML applications.
How might we develop reliable LtD and xAI methods when experts may disagree on the same deferred instance or explanation? These questions are often ignored when building modern decision flows (see, e.g., Cabitza, Campagner, and Basile (2023); Fahimi et al. (2024); Srivastava, Heidari, and Krause (2019)). Importantly, addressing this gap requires a formalization of perception that aligns with the standard ML problem formulation, which is probabilistic (Pearl 1988, 2009; Goodfellow et al. 2016). In this work, we provide a first approach to this objective.
@@ 1.1 Our Contributions
How can we formally represent individuals who disagree in their interpretation of the same information? We first propose a probabilistic framework for perception in which individuals construct disagreeing probability distributions, $P(\mathbf{X})$, for the same information $\mathbf{X}$. Under this framework, for example, we can explain why one expert assigns a higher probability to a positive classification for a deferred instance compared to another expert. We then propose causal perception by extending the initial formulation to account for probabilistic causal reasoning. Here, individuals are equipped with structural causal models (SCMs) (Pearl 2009) that not only describe $P(\mathbf{X})$ but also govern how they reason about hypothetical scenarios. As a result, individuals may disagree not only on observed distributions but also on interventional and counterfactual ones. Under this framework, we can describe how the two experts arrive at different classification probabilities by answering counterfactual questions, e.g., on a deferred instance. Our framework draws on foundational work in probabilistic and causal reasoning under uncertainty (Pearl 1988, 2009). We adopt a particular view of perception as disagreement in the interpretation of information conveyed through probabilities.
We define two kinds of causal perception: structural and parametrical. In structural perception individuals disagree on the cause-effect relationships, while in parametrical perception individual disagree on the causal effect between an established cause-effect pair. We complement these two kinds of causal perception by providing an initial approach to model the processes by which individuals construct their SCM. Further, we discuss areas of interest within fair ML and point at promising directions of future work involving perception, such as situated bias and reformulating sensitive attributes as loaded attributes. To showcase our framework, we present a series of examples involving ML applications and individual experts.
In the rest of the paper, we introduce the necessary background in Section 2. We then define probabilistic perception, or perception, in Section 3 and causal probabilistic perception, or causal perception, in Section 4. We present a first approach to the construction of the individual SCMs behind causal perception in Section 5. We discuss the fair ML implications in Section 6. We conclude in Section 7.
@@ 1.2 Related Work
Perception has been studied mainly by psychologist, with views on it varying within the field. We view perception as individual differences in interpreting the same information shaped by mental heuristics and expressed as probabilities; we discuss our choice later in Section 3. The focus within such view has been on the representativeness heuristic, in which an event is made to be more representative of a class than what it actually is, as measured by a higher probability (Tversky and Kahneman 1974, 1981, 1983). Bayesian modeling—in which the baseline representativeness of an event (the prior) is adjusted by the agent based on her experience (the posterior)—remains the common approach for modeling explicitly the representativeness heuristic and, thus, implicitly perception (Costello 2009; Bordalo et al. 2016; Tentori 2022). Such view belongs to a cognitive-bias-driven interpretation of perception in which perception influences rational decision-making under uncertainty (Kahneman 2011). There are, for instance, constructivist (Gregory 1970) and sensorial (Goldstein 1996) views that study the interpretative aspect of perception. We are the first to use SCMs under this view of perception, which allows to frame it also in terms of hypothetical distributions that represent counterfactual reasoning. Further, we extend previous works since SCMs, or causal Bayesian networks (Pearl 2009), include Bayesian reasoning. The proposed causal perception framework, thus, can be viewed as a comparison of two individuals’ posterior judgments and the causal theories that underlie those judgments.
@@
ML researchers are increasingly interested in cognitive biases (see, e.g., Bengio (2019); Booch et al. (2021)). Such works focus on how to create ML-driven intelligent systems that improve over and potentially replace the irrational human expert. It is an active and interdisciplinary line of research (see, e.g., Agrawal, Gans, and Goldfarb (2019)). Our work adds to this growing line of research by making use of causal reasoning as a basis for modeling human-like reasoning (Schölkopf 2022; Pearl 2009). Further, there are narrower lines of ML work that study these cognitive biases and their impact on the human user of the ML application. We highlight those studying the human-in-the-loop problem from a fairness and accuracy perspective (see, e.g., De et al. (2020); Mozannar et al. (2023); Palomba et al. (2025)), which includes LtD (Madras, Pitassi, and Zemel 2018) and xAI (Rong et al. 2024). Such works consider the human’s interaction with the ML model, incorporating it within the problem formulation. With some exceptions (Caraban and Karapanos 2020; Rastogi et al. 2022; Yang, Folke, and Shafto 2025), these works, however, do not give agency to the human user and formulate it as an additional and costly decision-maker, ignoring any influence from individual experience. Different from these works, we explore the setting of having multiple users that disagree on the information provided. In that sense, our work gives more agency to the user by recognizing the role of perception among a set of heterogeneous users.
@@
The role of these cognitive biases in fair ML is largely unexplored, with some exceptions. Some of these works show how the cognitive biases can be exacerbated by the xAI techniques used by the expert to explain a model (Bertrand et al. 2022) and can affect the expert’s evaluation of a model’s output (Echterhoff, Yarmand, and McAuley 2022). Further, given the problem of context-aware fairness, where we recognize that fairness has different meanings across humans, works like Srivastava, Heidari, and Krause (2019) and Yaghini, Krause, and Heidari (2021) design user experiments to test for the human perception of fairness. These experiments show that the fairness of an outcome can be judged differently depending on who is the individual judging. We add to these works by formalizing perception itself through causal reasoning. In Section 5.3 we further discuss perception within the causal fair ML literature.
@@ 1.3 A Running Example
Throughout the paper we use a series of ML-enabled decision flow scenarios based on Example 1.
@@
(College Admissions) An admissions officer (the decision maker or DM) chooses the incoming class based on the applicants’ profiles. Assume a decision flow in which the officer admits, $Y=1$, or rejects, $Y=0$, applicants based on their SAT results, $X_{1}$, high-school GPA, $X_{2}$, and suitability scores $f(X_{1},X_{2})=G\in[0,1]$ where $f$ is a ML model trained by the college. The officer, through the applicants’ motivation letters, also has access to their address, $Z$. Consider three scenarios in which the officer relies on $f$ with varying degrees for the decision:
@@
The above example is an extended version of Kleinberg et al. (2019)’s college admissions tiebreaker example. Scenarios (i) and (ii) represent a partially automated decision flow in which $f$ aids the DM, while scenario (iii) represents a fully automated decision flow in which $f$ replaces the DM. These are high-level but common ML-enabled scenarios. Additional context is provided for each scenario as we extend Example 1 moving forward.
@@ 2 Background
In this work, we represent information as a set of $p$ discrete random variables $\mathbf{X}=X_{1},\dots,X_{p}$. Let $P(\mathbf{X})$ denote the joint probability distribution of $\mathbf{X}$ with $P(\mathbf{x})$ representing the joint probability that $\mathbf{X}$ equals the $p$ instances $\mathbf{x}=x_{1},\dots,x_{p}$. We treat individuals, like a human user, and objects, like a ML application, as agents.
@@
Borrowing from signaling games (Spence 1973), we consider two types of agents. A sender $S\in\mathcal{S}$, with $\mathcal{S}$ denoting the set of senders, is an agent that provides information while a receiver $R\in\mathcal{R}$, with $\mathcal{R}$ denoting the set of receivers, is an agent that interprets the provided information. We only borrow this distinction as our agents are non-strategic and engage in a single one-shot game, meaning that the sender sends and the receiver receives information once without tricking one another. For an overview of signaling games, see Sobel (2020).
@@
We focus on the additive noise models (ANMs), where $f_{j}$ is a linear transformation:
@@
with $\beta_{ij}\in\mathbb{R}$ denoting the causal weight of the $i$-th parent $pa(j)_{i}$ of $X_{j}$. Our discussion applies to all SCMs, but we consider the ANM case for readability.
@@
which simplifies reasoning about $P(\mathbf{X})$, as it states that $X_{i}$ is conditionally independent of its non-descendants given its parents $X_{pa(i)}$. This property is known as the Markovian condition (Peters, Janzing, and Schölkopf 2017).
For the hypothetical scenarios, or what if, it is possible to generate new distributions of $P(\mathbf{X})$ by intervening the SCM $\mathcal{M}$. This is because SCMs build on an interventionist account of causality (Woodward 2005). An intervention on a single variable $X_{i}$ is done via the do-operator, $do(X_{i}=x_{i})$, which replaces the structural equation in $\mathbf{F}$ for the variable $X_{i}$ with the value $x_{i}$. Interventions apply similarly for multiple variables, $do(X_{i}=x_{i},X_{j}=x_{j})$, replacing the structural equations for each variable individually. Let $\mathcal{I}_{\mathbf{X}}$ denote the set of all interventions, which is an index set with each index representing a specific intervention on the variables $\mathbf{X}$. We use $\emptyset\in\mathcal{I}_{\mathbf{X}}$ to denote the null intervention. As Rubenstein et al. (2017) demonstrate, $\mathcal{I}_{\mathbf{X}}$ has a natural partial ordering, in which for interventions $i,j\in\mathcal{I}_{\mathbf{X}},\,i\leq_{\mathbf{X}}j$ if and only if $i$ intervenes on a subset of the variables that $j$ intervenes on and sets them equal to the same values as $j$. For instance, $do(X_{i}=x_{i})\leq_{\mathbf{X}}do(X_{i}=x_{i},X_{j}=x_{j})$. It means the $j$ intervention can be done after the $i$ intervention without needing to change the modifications done by the $i$ intervention on the SCM $\mathcal{M}$.
@@
By definition, $P(\mathbf{\mathbf{X}})\in\mathcal{P}_{\mathbf{X}}$. Further, $\mathcal{P}_{\mathbf{X}}$ is a singleton comprised of $P(\mathbf{X})$ when $\mathcal{I}_{\mathbf{X}}=\{\emptyset\}$. Furthermore, intuitively, $\mathcal{P}_{\mathbf{X}}$ represents all possible ways of reasoning causally about variables $\mathbf{X}$ as implied by a SCM $\mathcal{M}$.
@@ 3 Perception of Probability
We first present perception in its most basic form. The goal is to formulate when two individuals, the receivers, interpret differently the same information from a ML application, formalized as the sender. We consider, in particular, the setting in which these individuals act as some sort of decision maker within a decision flow.
We deal with perception of probability as we assume that the individuals manifest their interpretations of the information received using probabilities. This assumption is rooted in how humans reason under uncertainty, which has a long and established tradition across multiple fields that study bias in human decision-making (see, e.g., Kahneman (2011); Thaler and Sunstein (2008)). We find this to be a reasonable assumption. Whether it is because the setting is inherently stochastic or the individual cannot possibly observe all the variables involved, there is always some degree of uncertainty involved in decision flows (Pearl 1988). Probabilistic reasoning helps capture the uncertainty behind making a choice based on a ML application’s outcome. Further, probability theory is one of the pillars of ML (Goodfellow et al. 2016). By describing perception in terms of probabilities, we can relate it to standard ML problem formulations. Let us now offer a preliminary definition of perception of probability entirely based on this intuition.
@@
(Perception) For receivers $R_{i},R_{j}\in\mathcal{R}$, given the information by sender $S\in\mathcal{S}$ in the form of a random variable $X$, perception occurs when, for a threshold $\epsilon\in\mathbb{R}^{+}$, we have that:
$d\left(P_{R_{i}}(X),P_{R_{j}}(X)\right)>\epsilon$ (5)
@@
(From variable to variables) Definition 3.1 naturally applies to the $p$ random variables $\mathbf{X}$. In that case, $P_{R_{i}}(\mathbf{X})$ and $P_{R_{j}}(\mathbf{X})$ represent the joint probability distributions of the receivers for information $\mathbf{X}$. Further, we can consider the particular case in which one of the random variables denotes the target $Y$ and $P_{R_{i}}(Y|\mathbf{X})$ and $P_{R_{j}}(Y|\mathbf{X})$ represent the conditional joint probability distributions of the receivers for information $\mathbf{X}$ with p1p-1.
@@
(Instance-based perception) Implicit to Definition 3.1 is a disagreement at the instance level $X=x$, meaning $d\big{(}P_{R_{i}}(x),P_{R_{j}}(x)\big{)}>\epsilon$ where $P_{R_{i}}(x)$ and $P_{R_{i}}(x)$ are probabilities and $d(\cdot,\cdot)$ a suitable distance metric. [^2] Here, the focus is on the specific instance $x$ as, indeed, distance among probability distributions in Equation 5 may be $\leq\epsilon$, with still a number of instances having distance $\epsilon$. Intuitively, if a receiver can construct the probability distribution $P(X)$, then it can also compute the probability of any instance $x\in X$. It is also the case for the joint probabilities of the $p$ instances in $\mathbf{x}$ of the $p$ random variables in $\mathbf{X}$.
With Definition 3.1, we are interested in capturing how perception—as a disagreement in probability—would look like formally. We take for granted the process behind how these two receivers construct(ed) $P(X)$ and focus on the fact that there is a probabilistic disagreement between them. The fact that there is disagreement of probability between $R_{i}$ and $R_{j}$ only has an impact when these two receivers must make a decision within the same decision-making context based on $P(X)$. If $d\big{(}P_{R_{i}}(X),P_{R_{j}}(X)\big{)}\leq\epsilon$, then we would not have a disagreement and we would be indifferent between $R_{i}$ and $R_{j}$. In that case, these two receivers, for our purposes, would be indistinguishable within said context. Let us consider the next example.
@@
Let us consider Example 1, scenario (i). Suppose two admissions officers, $R_{1}$ and $R_{2}$, receive the same deferred applicant by the abstaining ML model $f$ with profile $\mathbf{x}=\langle x_{1},x_{2}\rangle$. Suppose that perception occurs as each officer constructs different $P(Y|X_{1},X_{2})$, such that $P_{R_{1}}(Y|x_{1},x_{2})>P_{R_{2}}(Y|x_{1},x_{2})$, for which officer $R_{1}$ accepts while $R_{2}$ rejects the applicant.
As Example 2 illustrates, Definition 3.1 dwells exclusively with the observational probability distribution or the what is. It does not allow for hypothetical scenarios, and the applicant is judged based on what is observed in his or her profile. There is no, for instance, counterfactual reasoning involved when deciding over the deferred applicant. We expand such a setting in Section 4.